OpenStudy (sloppycanada):

From the window 20 feet above the ground, the angle of the elevation to the the top of a building across the street is 78 degrees, and the angle of depression to the bse of the same building is 15 degrees. Find the height of the building across the street.

OpenStudy (sloppycanada):

@Shalante

OpenStudy (koikkara):

 Then, The height of the building: \(h = 20 + (20/tan15)*tan78 \) h = 20 + 351.16 h = 371.16 ft @sloppycanada \(\color{blue}{\huge\tt{Nice~}}\) \(\huge\mathcal{to~meet~}\) \(\color{red}{\huge\tt{You~!}}\)